To determine an athlete's body fat, she is weighed first in air and then again while she's completely underwater. It is found that she weighs 690 N when weighed in air and 40.0 N when weighed underwater. What is her average density? 
Answer #1

mass of the her, m = 690/9.8

= 70.4 kg

Let V is the volume of her body.


from the given data, buoynat force acting on the body, B = 690 - 40

= 650 N

we know, Buoynat force = weight of the dispalced water

B = rho_water*V*g

V = B/(rho_water*g)

= 650/(1000*9.8)

= 0.0663 m^3

so, average density of her, density = mass/volume

= 70.4/0.0663

= 1061.4 kg/m^3